# Leave No Document Behind: Benchmarking Long-Context LLMs with Extended Multi-Doc QA

### v1

@@ Abstract
Long-context modeling capabilities have garnered widespread attention, leading to the emergence of Large Language Models (LLMs) with ultra-context windows. Meanwhile, benchmarks for evaluating long-context LLMs are gradually catching up. However, existing benchmarks employ irrelevant noise texts to artificially extend the length of test cases, diverging from the real-world scenarios of long-context applications. To bridge this gap, we propose a novel long-context benchmark, Loong, aligning with realistic scenarios through extended multi-document question answering (QA). Unlike typical document QA, in Loong’s test cases, each document is relevant to the final answer, ignoring any document will lead to the failure of the answer. Furthermore, Loong introduces four types of tasks with a range of context lengths: Spotlight Locating, Comparison, Clustering, and Chain of Reasoning, to facilitate a more realistic and comprehensive evaluation of long-context understanding. Extensive experiments indicate that existing long-context language models still exhibit considerable potential for enhancement. Retrieval augmented generation (RAG) achieves poor performance, demonstrating that Loong can reliably assess the model’s long-context modeling capabilities. [^1] The code and benchmark are available at https://github.com/MozerWang/Loong
@@
Minzheng Wang1,3thanks: Equal contribution., Longze Chen2,311footnotemark: 1, Cheng Fu4, Shengyi Liao4, Xinghua Zhang4, Bingli Wu4 Haiyang Yu4, Nan Xu1, Lei Zhang2,3, Run Luo2,3, Yunshui Li2,3, Min Yang2thanks: Corresponding author., Fei Huang4, Yongbin Li422footnotemark: 2 1MAIS, Institute of Automation, Chinese Academy of Sciences 2Shenzhen Institute of Advanced Technology, Chinese Academy of Sciences 3School of Artificial Intelligence, University of Chinese Academy of Sciences 4Alibaba Group 🖂: wangminzheng2023@ia.ac.cn;  lz.chen2@siat.ac.cn
@@ 1 Introduction
Large Language Models (LLMs) have exhibited remarkable proficiency in diverse downstream applications OpenAI (2023). Recent works focus on scaling up the context window of LLMs (Xiong et al., 2023; Peng et al., 2023; Chen et al., 2024b), which is crucial for LLMs in handling complex tasks that require delving deeply into long texts. A few LLM (e.g. GPT-4o, Gemini-Pro) websites have been equipped with the intelligent document analysis function, allowing users to upload documents for answering queries. Meanwhile, retrieval-augmented generation (RAG) has been a commonly used framework that prompts LLMs with multiple relevant retrieved contents and can significantly improve model performance Wu et al. (2024); Chen et al. (2024a). These demand the model leverage its long-context capability to conduct an in-depth analysis of multiple long documents.
@@
L-Eval (An et al., 2023)
LongBench (Bai et al., 2023b)
Marathon (Zhang et al., 2023)
LooGLE (Li et al., 2023)
InfiniteBench (Zhang et al., 2024)
@@
However, there remains a lack of appropriate benchmarks for evaluating long-context understanding in real-world multi-document scenarios. Multi-document input as long-context modeling possesses extensive application scenarios of LLMs, such as analysis of financial reports over the years. Nevertheless, most existing benchmarks only place emphasis on single-document long contexts (An et al., 2023; Li et al., 2023; Kamradt, 2023) or involve multi-document question answering settings by adding distracting information to the input of existing short-context QA datasets (Hsieh et al., 2024). As shown in Figure 1, evidence supporting the answer in previous benchmarks is relatively centralized, such as being contained within a single document. Yet, such a centralized distribution of evidence may cause the model to overlook certain documents and take shortcuts to formulate an answer, complicating the modeling of the real context length. Moreover, the prevalent evaluation tasks, such as “needle in a haystack” (NIAH) (Kamradt, 2023), only scratch the surface of long-context understanding by searching from context, far from real-world demands.
@@
We commence with “$\mathtt{leave}$ $\mathtt{no}$ $\mathtt{document}$ $\mathtt{behind}$” and scatter the evidence across multi-document long contexts. In this context, bypassing any document will lead to an erroneous answer, which better tests the long-context modeling ability. To this end, this paper develops Loong, an innovative benchmark crafted to evaluate the long-context ability of LLMs across multiple documents in real-world scenarios. Loong typically consists of 11 documents per test instance on average, spanning three real-world scenarios in English and Chinese: (1) Financial Reports, (2) Legal Cases, and (3) Academic Papers. Meanwhile, Loong introduces new evaluation tasks from the perspectives of Spotlight Locating, Comparison, Clustering, and Chain of Reasoning. Furthermore, Loong features inputs of varying lengths (e.g., 10K-50K, 50K-100K, 100K-200K, >200K) and evaluation tasks of diverse difficulty, enabling fine-grained assessment of LLMs across different context lengths and task complexities.
@@
We conduct extensive experiments on Loong to test the long-context modeling capabilities of several advanced LLMs. The empirical results show that even the current most powerful LLMs still struggle with the tasks in Loong, suggesting significant room for improvement in current LLMs. Furthermore, this paper conducts in-depth analyses regarding the behavior of long-context LLMs, involving RAG and the scaling law of context size. Our main contributions are summarized as follows:
@@
  • Loong primarily focuses on testing long-context ability of LLMs across multiple documents by scattering the evidence to examine the real length of long contexts.
@@
  • Loong offers evaluation sets with varying input lengths and evaluation tasks of differing difficulty, encompassing novel task categories as well as common multi-document scenarios.
@@
  • All test instances are newly annotated and checked to guarantee quality. Extensive experiments and analyses deeply unveil the long-context modeling abilities of LLMs.
@@
Figure 2: Showcase of four evaluation tasks in Loong (<$\mathtt{di}$>…</$\mathtt{di}$> marks the content of the i-th document). a) Spotlight Locating: Locate the evidences. b) Comparison: Locate and compare the evidences. c) Clustering: Locate and cluster the evidences into groups. d) Chain of Reasoning: Locate and reasoning along a logical chain.
@@ 2.1 Long-Context Language Models
Considering the quadratic complexity of Transformer (Vaswani et al., 2017), training LLMs with extensive context windows from scratch necessitates substantial computational resources, exceeding the capabilities of the general researchers. Consequently, recent studies have explored ways to expand the context length of these models during the fine-tuning stage. For example, PI (Chen et al., 2023), NTK-aware (bloc97, 2023), YaRN (Peng et al., 2023), Giraffe (Pal et al., 2023), Code LLaMA (Roziere et al., 2023), and PoSE (Zhu et al., 2023) adapts position embedding based on the rotary position encoding (RoPE) (Su et al., 2024), with only a few fine-tuning steps, the context length can be efficiently extended.
Another strong baseline for long-context modeling is the sliding window method. Various sliding window-based variants such as ALibi (Press et al., 2021), xPos (Sun et al., 2022), PCW (Ratner et al., 2022), LM-Infinit (Han et al., 2023), StreamingLLM (Xiao et al., 2023) are used to achieve efficient context scaling. Yet they diverge from the global perception characteristic of the Transformer, failing to exploit the entire context.
@@ 2.2 Long-Context Benchmarks
Longbench (Bai et al., 2023b), LooGLE  (Li et al., 2023) and Marathon (Zhang et al., 2023) are earlier benchmarks for comprehensive assessment of long context. However, the average length for most tasks is between 5k and 25k, far less than the window size of LCLMs. L-Eval (An et al., 2023), BAMBOO (Dong et al., 2023), CLongEval (Qiu et al., 2024) and InfiniteBench (Zhang et al., 2024) contain sufficiently long evaluation data, and the wide variety of tasks makes the assessment more comprehensive. RULER (Hsieh et al., 2024) creates a comprehensive testing method with flexibly adjustable length and difficulty, yet they only add distracting information to the input of existing short-context QA datasets.
While these long-context benchmarks have their own advantages, we still lack a benchmark that is sufficiently long, free from data contamination (Golchin and Surdeanu, 2023), and fully aligned with the real-world multi-document question answering scenario.
@@ 2.3 Retrieval Augmented Language Models
Leveraging long documents as external knowledge, Retrieval Augmented Language Models (RALMs) has achieved comparable or even better performance than LCLMs fine-tuned for specific tasks with long document. In previous study, RALMs could directly utilize the content retrieved during the inference phase. REPLUG (Shi et al., 2023) treats the language model as a black box and the retrieval component as an adjustable plug-and-play module. RETRO (Borgeaud et al., 2022) use a chunked cross-attention module to incorporate the retrieved text. Additionally, Xu et al. (2023) explored whether RALMs or LCLMs are more suitable for long-context tasks under a larger parameter setting. However, there is currently a lack of analysis on what tasks RALMs and LCLMs each excel at, thus making it difficult to determine which type a black box model belongs to.
@@ 3.1 Overview
The Loong benchmark comprises tasks across four categories: Spotlight Locating, Comparison, Clustering, and Chain of reasoning. To align with realistic scenarios, we collect documents from three domains: financial reports, academic papers, and legal cases. Furthermore, all tasks are presented in the question-answering format, which are all newly annotated by GPT-4o and humans. Totally, Loong includes 1600 test instances in both Chinese and English, featuring four sets with different intervals of context size: $\mathtt{Set1}$ (10-50K), $\mathtt{Set2}$ (50-100K), $\mathtt{Set3}$ (100-200K) and $\mathtt{Set4}$ (200-250K). We use tiktoken [^2] https://platform.openai.com/tokenizer tokenizer to tokenize the input and report the number of tokens. Table 2 and Appendix C show the details of data statistics. The following sections will provide a detailed description of the evaluation task and benchmark construction.
@@ 3.2 Evaluation Task
Based on various multi-document semantic relationships and LLMs’ handling of multi-document input, we propose new task categories for multi-document long-context modeling and closer alignment with real-world scenarios. Figure 2 illustrates the evaluation tasks of the Loong benchmark. Appendix B shows the detailed test case and prompt of each task. Appendix E presents a comparison of evidence distribution between Loong and LongBench.
@@ 3.2.1 Spotlight Locating
The spotlight locating task is designed to assess the model’s capability for knowledge localization, which constitutes the foundation ability of long-context processing. In this task, the evidences are contained in only one of multiple documents, which is the atomic setting of the key information locating. Spotlight locating task is aimed at examining the LLMs’ ability to search the evidence within one document from multiple ones. Other documents, which are in the same domain and have similar semantics as the document but are unrelated to the question, will serve as noise texts. The upper left of Figure 2 provides an example of the spotlight locating task.
@@ 3.2.3 Clustering
The clustering task entails an assessment of the model’s ability to cluster key information based on specific conditions across multi-document long contexts. This task claims that LLMs cluster relevant evidence scattered in multiple documents based on the specified criteria. Furthermore, it necessitates the extraction of pertinent information from documents and the integration of this information by grouping according to conditions.
@@ 3.3.1 Data Collection
We established six criteria for the manual collection of the required English and Chinese documents: (1) Timeliness: The majority of the documents are the latest ones from the year 2024; (2) Accessibility: The data is publicly available and permitted for download and collection; (3) Appropriate Length: Collecting longer documents as much as possible and ensure they fit within the four designated length sets; (4) Parseability: Chosen documents are easy to process and parse, facilitating conversion into natural language text; (5) Categorizability: Documents can be manually sorted based on certain attributes, such as case type, research theme, or company category, allowing for organized archival; (6) Authoritativeness: All documents are collected from scratch from official websites (e.g. China Judge Online [^3] https://wenshu.court.gov.cn/, U.S. SEC [^4] https://www.sec.gov/, cninf [^5] http://www.cninfo.com.cn/, Arxiv [^6] https://arxiv.org/, Semantic Scholar [^7] https://www.semanticscholar.org/), ensuring the quality and authority of the documents.
@@ 4.1 Experimental Setup
Models We evaluate six advanced long-context LLMs, with their context window sizes ranging from 128K to 1000K, including API-based LLMs: GPT-4o-128K (OpenAI, 2023), Gemini-Pro1.5-1000K (Reid et al., 2024), Claude3.5-Sonnet-200K (Anthropic, 2024b), Claude3-Haiku-200K (Anthropic, 2024a), Kimi-Chat-200K [^8] https://kimi.moonshot.cn/ and Open-sourced LLMs: Qwen2-72B-Instruct-128K (Bai et al., 2023a), GLM4-9B-Chat-1000K (Du et al., 2022).
Evaluation Metric In the long-context question-answering scenarios, traditional evaluation metrics F1 and Rouge-L may lead to inaccurate responses. Recent research (Liu et al., 2024; Wang et al., 2024) indicates that the GPT-4 (OpenAI, 2023) evaluator demonstrates high consistency with human evaluations, making it a reasonably reliable annotator. Building on these considerations, we prompt GPT-4 as a judge to evaluate the model’s output based on the golden answer and the question’s requirements from three aspects: Accuracy, Hallucinations, and Completeness, scoring from 0 to 100. For a detailed prompt, please refer to the Appendix A. We also design two indicators: (1) Avg Scores: the average value of scores given by GPT-4 for all questions; (2) Perfect Rate: the proportion of cases scoring 100 out of the total cases. The latter is a more stringent evaluation metric compared to the former.

### v2

@@ Abstract
Long-context modeling capabilities have garnered widespread attention, leading to the emergence of Large Language Models (LLMs) with ultra-context windows. Meanwhile, benchmarks for evaluating long-context LLMs are gradually catching up. However, existing benchmarks employ irrelevant noise texts to artificially extend the length of test cases, diverging from the real-world scenarios of long-context applications. To bridge this gap, we propose a novel long-context benchmark, Loong, aligning with realistic scenarios through extended multi-document question answering (QA). Unlike typical document QA, in Loong’s test cases, each document is relevant to the final answer, ignoring any document will lead to the failure of the answer. Furthermore, Loong introduces four types of tasks with a range of context lengths: Spotlight Locating, Comparison, Clustering, and Chain of Reasoning, to facilitate a more realistic and comprehensive evaluation of long-context understanding. Extensive experiments indicate that existing long-context language models still exhibit considerable potential for enhancement. Retrieval augmented generation (RAG) achieves poor performance, demonstrating that Loong can reliably assess the model’s long-context modeling capabilities. [^1] The code and benchmark are available at https://github.com/MozerWang/Loong and https://github.com/AlibabaResearch/DAMO-ConvAI/tree/main/Loong
@@
Minzheng Wang1,3thanks: Equal contribution., Longze Chen2,311footnotemark: 1, Cheng Fu4, Shengyi Liao4, Xinghua Zhang4 Bingli Wu4, Haiyang Yu4, Nan Xu1, Lei Zhang2,3, Run Luo2,3, Yunshui Li2,3 Min Yang2thanks: Corresponding author., Fei Huang4, Yongbin Li422footnotemark: 2 1 MAIS, Institute of Automation, Chinese Academy of Sciences 2 Shenzhen Institute of Advanced Technology, Chinese Academy of Sciences 3 School of Artificial Intelligence, University of Chinese Academy of Sciences 4 Alibaba Group 🖂: wangminzheng2023@ia.ac.cn;  lz.chen2@siat.ac.cn
@@ 1 Introduction
Large Language Models (LLMs) have exhibited remarkable proficiency in diverse downstream applications OpenAI (2023). Recent works focus on scaling up the context window of LLMs (Xiong et al., 2024; Peng et al., 2024; Chen et al., 2024b), which is crucial for LLMs in handling complex tasks that require delving deeply into long texts. A few LLM (e.g. GPT-4o, Gemini-Pro) websites have been equipped with the intelligent document analysis function, allowing users to upload documents for answering queries. These demand the model leverage its long-context capability to conduct an in-depth analysis of multiple long documents. Meanwhile, retrieval-augmented generation (RAG) has emerged as a widely adopted framework that prompts LLMs with multiple relevant retrieved contents and can significantly improve model performance Wu et al. (2024); Chen et al. (2024a).
@@
L-Eval (An et al., 2024)
LongBench (Bai et al., 2024)
Marathon (Zhang et al., 2024a)
LooGLE (Li et al., 2024)
InfiniteBench (Zhang et al., 2024b)
@@
However, there remains a lack of appropriate benchmarks for evaluating long-context understanding in real-world multi-document scenarios. Multi-document input as long-context modeling possesses extensive application scenarios of LLMs, such as analysis of financial reports over the years. Nevertheless, most existing benchmarks only place emphasis on single-document long contexts (An et al., 2024; Li et al., 2024; Kamradt, 2023) or involve multi-document question answering settings by adding distracting information to the input of existing short-context QA datasets (Hsieh et al., 2024). As shown in Figure 1, evidence supporting the answer in previous benchmarks is relatively centralized, such as being contained within a single document. Yet, such a centralized distribution of evidence may cause the model to overlook certain documents and take shortcuts to formulate an answer, simplifying the modeling of the real context. Moreover, the prevalent evaluation tasks, such as “needle in a haystack” (NIAH) (Kamradt, 2023), only scratch the surface of long-context understanding by searching from context, far from real-world demands.
@@
We commence with “$\mathtt{leave}$ $\mathtt{no}$ $\mathtt{document}$ $\mathtt{behind}$” and scatter the evidence across multi-document long contexts. In this context, bypassing any document will lead to an erroneous answer, which better tests the long-context modeling ability. To this end, this paper develops Loong, an innovative benchmark crafted to evaluate the long-context ability of LLMs across multiple documents in real-world scenarios. Loong typically consists of 11 documents per test instance on average, nearly all of which are selected from the year 2024, spanning three real-world scenarios in English and Chinese: (1) Financial Reports, (2) Legal Cases, and (3) Academic Papers. Meanwhile, Loong introduces new evaluation tasks from the perspectives of Spotlight Locating, Comparison, Clustering, and Chain of Reasoning, every test case is newly annotated. Furthermore, Loong features inputs of varying lengths (e.g., 10K-50K, 50K-100K, 100K-200K, >200K) and evaluation tasks of diverse difficulty, enabling fine-grained assessment of LLMs across different context lengths and task complexities.
@@
We conduct extensive experiments on Loong to test the long-context modeling capabilities of several advanced LLMs. The empirical results show that even the current most powerful LLMs still struggle with the tasks in Loong, suggesting significant room for improvement in current LLMs. Furthermore, this paper conducts in-depth analyses regarding the behavior of long-context LLMs involving RAG and the scaling law of context size. Our main contributions are summarized as follows:
@@
  • Loong primarily focuses on testing the long-context modeling ability of LLMs across multiple documents by scattering the evidence to examine the actual length of long contexts.
@@
  • Loong offers evaluation sets with varying input lengths and evaluation tasks of differing difficulty, encompassing novel task categories. All test instances are newly annotated and checked to guarantee quality.
@@
  • Extensive experiments and analyses deeply unveil the performance of Long-Context LLMs, also offering insights into improving Long-Context modeling abilities.
@@
Figure 2: Showcase of four evaluation tasks in Loong (<$\mathtt{di}$>…</$\mathtt{di}$> marks the content of the i-th document). a) Spotlight Locating: Locate the evidence. b) Comparison: Locate and compare the evidence. c) Clustering: Locate and cluster the evidence into groups. d) Chain of Reasoning: Locate and reasoning along a logical chain.
@@ 2.1 Long-Context Language Models
Considering the quadratic complexity of Transformer (Vaswani et al., 2017), training LLMs with extensive context windows from scratch necessitates substantial computational resources, exceeding the capabilities of the general researchers. Consequently, recent studies have explored ways to expand the context length of these models during the fine-tuning stage. For example, PI (Chen et al., 2023), NTK-aware (bloc97, 2023), YaRN (Peng et al., 2024), Giraffe (Pal et al., 2023), Code LLaMA (Roziere et al., 2023), and PoSE (Zhu et al., 2024) adapts position embedding based on the rotary position encoding (RoPE) (Su et al., 2024), with only a few fine-tuning steps, the context length can be efficiently extended.
Another strong baseline for long-context modeling is the sliding window method. Various sliding window-based variants such as ALibi (Press et al., 2022), xPos (Sun et al., 2023), PCW (Ratner et al., 2023), LM-Infinit (Han et al., 2024), StreamingLLM (Xiao et al., 2023) are used to achieve efficient context scaling. Yet they diverge from the global perception characteristic of the Transformer, failing to exploit the entire context.
@@ 2.2 Long-Context Benchmarks
Longbench (Bai et al., 2024), LooGLE  (Li et al., 2024) and Marathon (Zhang et al., 2024a) are earlier benchmarks for comprehensive assessment of long context. However, the average length for most tasks is between 5k and 25k, far less than the window size of LCLMs. L-Eval (An et al., 2024), BAMBOO (Dong et al., 2024), CLongEval (Qiu et al., 2024) and InfiniteBench (Zhang et al., 2024b) contain sufficiently long evaluation data, and the wide variety of tasks makes the assessment more comprehensive. RULER (Hsieh et al., 2024) creates a comprehensive testing method with flexibly adjustable length and difficulty, yet they only add distracting information to the input of existing short-context QA datasets.
While these long-context benchmarks have their advantages, there still lack a benchmark that is sufficiently long, free from data contamination (Golchin and Surdeanu, 2023), and fully aligned with the real-world multi-document question answering scenario. We conduct a detailed comparison with existing works in Table 1.
@@ 2.3 Retrieval Augmented Language Models
Leveraging long documents as external knowledge, Retrieval Augmented Language Models (RALMs) have achieved comparable or even better performance than LCLMs fine-tuned for specific tasks with long documents. In previous study, RALMs could directly utilize the content retrieved during the inference phase. REPLUG (Shi et al., 2024) treats the language model as a black box and the retrieval component as an adjustable plug-and-play module. RETRO (Borgeaud et al., 2022) use a chunked cross-attention module to incorporate the retrieved text. Additionally, Xu et al. (2024) explored whether RALMs or LCLMs are more suitable for long-context tasks under a larger parameter setting. However, there is currently a lack of analysis on what tasks RALMs and LCLMs each excel at, thus making it difficult to determine which type a black box model belongs to.
@@ 3.1 Overview
The Loong benchmark comprises tasks across four categories: Spotlight Locating, Comparison, Clustering, and Chain of reasoning. To align with realistic scenarios, we collect documents from three domains: financial reports, academic papers, and legal cases. Furthermore, all tasks are presented in the question-answering format, which are all newly annotated by GPT-4o and humans. Totally, Loong includes 1600 test instances in both Chinese and English, featuring four sets with different intervals of context size: $\mathtt{Set1}$ (10-50K), $\mathtt{Set2}$ (50-100K), $\mathtt{Set3}$ (100-200K) and $\mathtt{Set4}$ (200-250K). We use tiktoken [^2] https://platform.openai.com/tokenizer tokenizer to tokenize the input and report the number of tokens. Table 2 and Appendix C show the details of data statistics. Appendix E presents a comparison of evidence distribution between Loong and LongBench. The following sections will provide a detailed description of the evaluation task and benchmark construction.
@@ 3.2 Evaluation Task
Based on various multi-document semantic relationships and LLMs’ handling of multi-document input, we propose new task categories for multi-document long-context modeling and closer alignment with real-world scenarios. Figure 2 illustrates the evaluation tasks of the Loong benchmark. Appendix B shows the detailed test case and prompt of each task.
@@ 3.2.1 Spotlight Locating
The spotlight locating task is designed to assess the model’s capability for knowledge localization, which constitutes the foundation ability of long-context processing. In this task, the evidence is contained in only one of multiple documents, which is the atomic setting of the key information locating. The spotlight locating task is aimed at examining the LLMs’ ability to search for evidence within one document from multiple ones. Other documents, which are in the same domain and have similar semantics as the document but are unrelated to the question, will serve as noise texts. The upper left of Figure 2 provides an example of the spotlight locating task.
@@ 3.2.3 Clustering
The clustering task entails an assessment of the model’s ability to aggregate multi-source information from multi-document long contexts. The task requires the LLM to perform two main functions: first, to extract relevant data from multiple documents, and second, to integrate this information based on specific criteria. Ultimately, the LLM must cluster related evidence spread across these documents into coherent groups.
@@ 3.3.1 Data Collection
We established six criteria for the manual collection of the required English and Chinese documents: (1) Timeliness: The majority of the documents are the latest ones from the year 2024; (2) Accessibility: The data is publicly available and permitted for download and collection; (3) Appropriate Length: Collecting longer documents as much as possible and ensure they fit within the four designated length sets; (4) Parseability: Chosen documents are easy to process and parse, facilitating conversion into natural language text; (5) Categorizability: Documents can be manually sorted based on certain attributes, such as case type, research theme, or company category, allowing for organized archival; (6) Authoritativeness: All documents are collected from scratch from official websites (e.g. China Judge Online, U.S. SEC, cninf, Arxiv, Semantic Scholar), ensuring the quality and authority of the documents. The detailed URL can be seen in Appendix H.
@@ 4.1 Experimental Setup
Models We evaluate six advanced long-context LLMs, with their context window sizes ranging from 128K to 1000K, including API-based LLMs: GPT-4o-128K (OpenAI, 2023), Gemini-1.5-pro-1000K (Reid et al., 2024), Claude3.5-Sonnet-200K (Anthropic, 2024b), Claude3-Haiku-200K (Anthropic, 2024a), Kimi-Chat-200K [^3] https://kimi.moonshot.cn/ and Open-sourced LLMs: Qwen2-72B-Instruct-128K (Bai et al., 2023), GLM4-9B-Chat-1000K (Du et al., 2022).
Evaluation Metric In the long-context question-answering scenarios, traditional evaluation metrics F1 and Rouge-L may lead to inaccurate responses. Recent research (Zhang et al., 2023; Liu et al., 2024; Wang et al., 2024) indicates that the GPT-4 (OpenAI, 2023) evaluator demonstrates high consistency with human evaluations, making it a reasonably reliable annotator. Building on these considerations, we prompt GPT-4 as a judge to evaluate the model’s output based on the golden answer and the question’s requirements from three aspects: Accuracy, Hallucinations, and Completeness, scoring from 0 to 100. For a detailed prompt, please refer to the Appendix A. We also design two indicators: (1) Avg Scores: the average value of scores given by GPT-4 for all questions; (2) Perfect Rate: the proportion of cases scoring 100 out of the total cases. The latter is a more stringent evaluation metric compared to the former.